OpenStudy (anonymous):

Which is an equation in slope-intercept form for the line passing through (5, 1) and (–4, 7) ?? A. 2/3x + 13/3 B. -2/3x - 13/3 C. -2/3x + 13/3 D. 2/3x - 13/3 help ??

OpenStudy (anonymous):

slope intercept form is y=mx+b m is slope, b is y-intercept. first, let's find m using two points. m = (y₂-y₁)/(x₂-x₁) = (7 - 1)/(-4 - 5) = 6 / -9 = -2/3 now, y = -2/3x + b to find b, put (5, 1) or (-4, 7) into the equation above. let's put (5, 1) into y = -2/3x + b 1 = -2/3*5 + b 1 = -10/3 + b 1 + 10/3 = -10/3 + b - 10/3 13/3 = b the answer is, y = -2/3x + 13/3 slope is -2/3, y-intercept is 13/3. if you are still confused, i want you to follow the link below that explains the concept of slope intercept form clearly. http://www.brightstorm.com/math/algebra/...

OpenStudy (anonymous):

Use the formula \[m=\frac{y_2-y_1}{x_2-x_1}\] to find the slope, and then use the coordinates of one of the point to find y-intercept (b).

OpenStudy (anonymous):

okay , thank you !!